# Efficient Conditional Gradient Methods for Solving Stochastic Convex Bilevel Optimization Problems

### v1

@@ Abstract
We propose efficient methods for solving stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic optimization problem. Existing methods often rely on costly projection or linear optimization oracles over complex sets, which limits scalability. To overcome this, we propose an iteratively regularized conditional gradient framework that leverages efficient linear optimization oracles exclusively over the base feasible set. Our proposed methods employ a vanishing regularization sequence that progressively emphasizes the inner problem while biasing towards desirable minimal outer objective solutions. Under standard convexity assumptions, we establish non-asymptotic convergence rates of $O(t^{-({1}/{2}-p)})$ for the outer objective and $O(t^{-p})$ for the inner objective, where $p\in(0,1/2)$ controls the regularization decay, in the one-sample stochastic setting, and $O(t^{-(1-p)})$ and $O(t^{-p})$ in the finite-sum setting using a mini-batch scheme, where $p\in(0,1)$. Experimental results on over-parametrized regression and $\ell_{1}$-constrained logistics regression tasks demonstrate the practical advantages of our approach over existing methods, confirming our theoretical findings.
@@ 1 Introduction
We consider stochastic convex bilevel optimization problems, where the goal is to minimize an outer stochastic convex objective function subject to the solution set of an inner stochastic convex optimization problem. Formally, the problem is defined as
$F_{\operatorname{opt}}:=\min_{x\in{\mathbb{R}}^{d}}\left\{F(x):={\mathbb{E}}[f(x,\theta)]:x\in X_{\operatorname{opt}}\right\},\ \text{where }X_{\operatorname{opt}}:=\operatorname*{arg\,min}_{z\in X}G(z):={\mathbb{E}}[g(z,\xi)].$ (1)
Here $X\subset\mathbb{R}^{d}$ is a convex and compact set, and both $F$ and $G$ are smooth and convex, making the optimization problem (1) convex. This framework has broad applicability, including hyper-parameter optimization [15, 16, 38], meta-learning [16, 42], reinforcement learning [31] and game theory [50]. In particular, problem (1) arises in learning applications where $G$ is the prediction error of a model $x$ on data, and $F$ is an auxiliary objective (e.g., regularization, or a validation data set). The dependence on data is through the expectation operators and the variables $\theta,\xi$. The bilevel formulation provides a tuning-free alternative to the usual regularization approach where we optimize $\sigma F(x)+G(x)$ for a carefully tuned parameter $\sigma$.
Despite its wide applications in machine learning there are three primary challenges in solving (1). First, we do not have an explicit representation of the optimal set $X_{\operatorname{opt}}$ in general, which prevents us from using some common operations in optimization such as projection onto or linear optimization over $X_{\operatorname{opt}}$. An alternative is to reformulate (1) using its value function formulation
@@ 1.1 Related works
Simple bilevel optimization extends Tikhonov regularization [52] beyond $\ell_{2}$ penalties by considering a blended objective $\sigma F(x)+G(x)$, where $\sigma>0$ modulates the trade-off between inner and outer objectives. As $\sigma\to 0$, the problem converges to a bilevel form that prioritizes the inner objective. Friedlander and Tseng [17] established fundamental conditions ensuring solution existence for $\sigma>0$, forming the basis for iterative regularization, where $\sigma$ is gradually reduced. This idea was first implemented by Cabot [3] in the unconstrained case $X={\mathbb{R}}^{d}$ using a decreasing sequence $\sigma_{t}$, and later extended to general convex constraints by Dutta and Pandit [13]. While these methods use proximal steps, which can be computationally expensive, Solodov [49] proposed a more efficient gradient-based alternative. Helou and Simões [23] further generalized the framework to non-smooth settings via an $\epsilon$-subgradient method. These works show asymptotic convergence but lack convergence rates. Amini and Yousefian [1] analyzed an iterative regularized projected gradient method and established a convergence rate of $O(t^{-(1/2-b)})$ for the inner objective, assuming non-smooth $F$ and $G$, compact $X$, and strong convexity of $f$. Kaushik and Yousefian [29] later removed the strong convexity assumption and proved rates of $O(t^{-b})$ and $O(t^{-(1/2-b)})$ for the inner and outer objectives, respectively, for any $b\in(0,1/2)$. Malitsky [34] studied an accelerated variant based on Tseng’s method [53], achieving an $o(t^{-1})$ rate for the inner problem. More recently, Merchav et al. [36] proposed an accelerated proximal scheme with a $O(\log(t+1)/t)$ rate under smoothness assumption and $O(t^{-2})$ / $O(1/t^{2-\gamma})$ under a Hölderian error bound condition. Our work also follows an iterative regularization approach, but avoids expensive projections or proximal operations by leveraging linear optimization oracles.
@@
Several important applications have domains $X$ where projection-based operations are inefficient, yet linear optimization over $X$ is easier. To address the limitations of projection-based methods, recent work has focused on projection-free bilevel optimization via linear optimization oracles. Jiang et al. [28] approximated $X_{\operatorname{opt}}$ by replacing the constraint $G(x)\leq G_{\operatorname{opt}}$ with a linear approximation, which removed the need for Slater’s condition, and applying the conditional gradient method, refining the approximation at each step. They established $O(t^{-1})$ convergence rates for both objectives under smoothness and compactness assumptions. However, their method requires a pre-specified tolerance parameter $\epsilon_{g}$, and only guarantees $(\epsilon_{g}/2)$-infeasibility. Doron and Shtern [11] proposed an alternative approach based on sublevel sets of the outer function $F$. Their method performs conditional gradient updates over sets of the form $X\times\{x:F(x)\leq\alpha\}$, with a surrogate $\hat{G}_{t}$ that is updated iteratively. They achieved rates of $O(t^{-1})$ for the inner and $O(t^{-1/2})$ for the outer objective under composite structure in $G$ and an error bound on $F$. While linear optimization over $X$ may be efficient, this is not always true for sets like $X\cap H$, where $H$ is a halfspace, or $\{F(x)\leq\alpha\}$ unless $F$ has a special structure. To address this, Giang-Tran et al. [19] introduced a conditional gradient-based iterative regularization scheme that only requires linear optimization over $X$, albeit with slower rates: $O(t^{-p})$ and $O(t^{-(1-p)})$ for the inner and outer problems, respectively, for any $p\in(0,1)$. We extend this framework to the stochastic setting, maintaining the same oracle-based reliance on linear optimization over $X$.
@@
Stochastic bilevel optimization remains less explored. Jalilzadeh et al. [26] developed an iterative regularization-based stochastic extragradient algorithm, requiring $O(1/\epsilon_{f}^{4})$ and $O(1/\epsilon_{g}^{4})$ stochastic gradient queries to achieve $\epsilon$-optimality for both objectives. Cao et al. [4] extended the projection-free framework of Jiang et al. [28] to the stochastic setting. Their methods achieve $\tilde{O}(t^{-1})$ convergence rate when the noise distributions have finite support, and $O(t^{-1/2})$ convergence rates under sub-Gaussian noise. However, their method requires linear optimization over intersections of $X$ with halfspaces. In contrast, our work relaxes this requirement by relying solely on linear optimization over the base set $X$.
@@
Several alternative approaches exist for simple bilevel optimization, including sequential averaging schemes [44, 46, 35], sublevel set-based methods [2, 5, 58, 20], accelerated algorithms [45, 54, 6, 5] and primal-dual strategies [47]. These methods are less closely related to our framework and do not address the stochastic setting. More general bilevel problems have seen considerable attention, particularly in Stackelberg game formulations where the upper and lower levels model leader-follower dynamics [18, 7, 30, 27, 9, 32, 56, 33]. Recent extensions include contextual bilevel optimization with exogenous variables [55, 24, 51] and problems with functional constraints at the lower level [29, 40]. In contrast, our focus is restricted to the stochastic convex formulation of simple bilevel problems.
@@ 1.2 Contributions
    One-sample stochastic setting: We develop a projection-free method for stochastic bilevel optimization that achieves high-probability convergence rates of $O(t^{-(1/2-p)}\sqrt{\log(dt^{2}/\delta)})$ for the outer-level problem and $O(t^{-p}\sqrt{\log(d/\delta)})$ for the inner-level problem, with probability at least $1-\delta$ for any $p\in(0,1/2)$. Compared to the state-of-the-art approach by Cao et al. [4, Algorithm 1], which achieves the $O(t^{-1/2}\sqrt{\log(Td/\delta)})$ rate for both the outer and inner problems, our method offers several key advantages. First, it eliminates the need for optimization over intersections of the base feasible set with halfspaces, requiring only linear optimization over the original constraint set. Second, it removes the dependency on their pre-specified tolerance parameter $\epsilon_{g}$, and the associated computationally expensive initialization procedure to find an $\epsilon_{g}$-optimal starting point. Thus, our algorithm enjoys anytime guarantees.
@@
    Finite-sum setting: For problems with $n$ component functions, we establish convergence rates of $O(t^{-(1-p)}\sqrt{\log(t^{2}/\delta)})$ (outer) and $O(t^{-p}\sqrt{\log(1/\delta)})$ (inner) with probability $1-\delta$, for any $p\in(0,1)$. Although Cao et al. [4, Algorithm 2] achieves improved $O(t^{-1}\log(T)\sqrt{\log(T/\delta)})$ rates for both objectives, it again relies on more computationally demanding halfspace-intersection oracles, requires an additional initialization step, and depends on a fixed stepsize.
@@
    Numerical Validation: We demonstrate the practical efficiency of our methods on over-parametrized regression and $\ell_{1}$-constrained logistic regression tasks, showing significant speedups over existing projection-based and projection-free approaches. The experiments validate both the convergence rates and computational advantages of our framework.
Proof of these results are provided in Appendix A. Additional results are presented in Appendix B.
2 Algorithms and convergence guarantees
In this section, we present two iterative regularization approaches for solving problem (1) in both stochastic and finite-sum settings. The core idea is to introduce a decreasing regularization sequence $\{\sigma_{t}\}_{t\geq 1}$, and at each iteration $t$, perform a gradient-based update on the composite objective $\sigma_{t}F+G$. As $\sigma_{t}$ decreases, the algorithm gradually shifts focus from $F$ to $G$, thereby steering $x_{t}$ toward points in $X_{\operatorname{opt}}$. At the same time the $\sigma_{t}F$ term encourages $x_{t}$ towards a point in $X_{\operatorname{opt}}$ that minimizes $F$. Giang-Tran et al. [19] applied this idea to design a projection-free algorithm for the deterministic setting, establishing convergence rates of $O(t^{-p})$ for the upper level and $O(t^{-(1-p)})$ for the lower level when using $\sigma_{t}=\varsigma(t+1)^{-p}$ with $p\in(0,1)$.
In the stochastic setting Eq. 1, however, computing exact gradients of the upper and lower objective functions is computationally expensive. To address this challenge, we rely on sample-based gradient estimators. Specifically, given mini-batch sample sets ${\cal S}_{f}:=\{\theta_{1},\ldots,\theta_{|{\cal S}_{f}|}\}$ and ${\cal S}_{g}:=\{\xi_{1},\ldots,\xi_{|{\cal S}_{g}|}\}$ drawn independently at each iteration, we define the empirical gradients as
$\nabla f_{{\cal S}_{f}}(x):=\frac{1}{|{\cal S}_{f}|}\sum_{\theta\in{\cal S}_{f}}\nabla f(x,\theta)\quad\&\quad\nabla g_{{\cal S}_{g}}(x):=\frac{1}{|{\cal S}_{g}|}\sum_{\xi\in{\cal S}_{g}}\nabla g(x,\xi),$
which provide unbiased estimates of $\nabla F(x)$ and $\nabla G(x)$, respectively. A naïve approach is to use these sample gradients directly, that is, $\widehat{\nabla F}(x)=\nabla f_{\mathcal{S}_{f}}(x)$ and $\widehat{\nabla G}(x)=\nabla g_{\mathcal{S}_{g}}(x)$. However, these estimators maintain constant variance over optimization time, which prevents achieving optimal convergence rates. To address this limitation, we implement two distinct variance-reduced gradient estimation schemes tailored to different problem settings.
  • $\diamond$
    One-sample stochastic setting: When we access single samples of $\theta$ and $\xi$ per iteration, we employ the STOchastic Recursive Momentum (STORM) estimator [8], which progressively reduces variance through momentum-based updates under sub-Gaussian noise assumptions. The resulting method, called the Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) algorithm, is summarized in Algorithm 1. We note that the use of the STORM estimator in projection-free methods was first introduced in [60].
  • $\diamond$
    Finite-sum setting: When the expectations in (1) are finite sums of size $n$, we use the the Stochastic Path-Integrated Differential EstimatoR (SPIDER[37, 14] with $\sqrt{n}$-sized mini-batches, achieving variance reduction through periodic gradient recomputation. The resulting method, called the Iteratively Regularized Finite-Sum Conditional Gradient (IR-FSCG) algorithm, is summarized in Algorithm 2. Here, for any $n\in{\mathbb{N}}$, $[n]:=\{1,\dots,n\}$. We note that the SPIDER estimator in projection-free methods has been previously used in [48, 57, 59, 22].
While a mini-batch scheme can be applied even when expectations are not finite sums, we emphasize that this approach only improves the constants in the convergence rates and enhances algorithmic stability without affecting the asymptotic rate. Additionally, if the upper- and lower-level problems have different numbers of component functions ($n_{f}\neq n_{g}$), Algorithm 2 can be easily modified by introducing $S_{f}$ and $q_{f}$ for the upper level, and $S_{g}$ and $q_{g}$ for the lower level.
Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$,$\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$.

### v2

# Conditional Gradient Methods with Standard LMO for Stochastic Simple Bilevel Optimization

@@ Abstract
We propose efficient methods for solving stochastic simple bilevel optimization problems with convex inner levels, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic optimization problem. Existing methods often rely on costly projection or linear optimization oracles over complex sets, limiting their scalability. To overcome this, we propose an iteratively regularized conditional gradient approach that leverages linear optimization oracles exclusively over the base feasible set. Our proposed methods employ a vanishing regularization sequence that progressively emphasizes the inner problem while biasing towards desirable minimal outer objective solutions. In the one-sample stochastic setting and under standard convexity assumptions, we establish non-asymptotic convergence rates of $\tilde{O}(t^{-1/4})$ for both the outer and inner objectives. In the finite-sum setting with a mini-batch scheme, the corresponding rates become $\tilde{O}(t^{-1/2})$. When the outer objective is nonconvex, we prove non-asymptotic convergence rates of $\tilde{O}(t^{-1/7})$ for both the outer and inner objectives in the one-sample stochastic setting, and $\tilde{O}(t^{-1/4})$ in the finite-sum setting. Experimental results on over-parametrized regression and dictionary learning tasks demonstrate the practical advantages of our approach over existing methods, confirming our theoretical findings.
@@ 1 Introduction
We consider stochastic simple bilevel optimization problems, where the goal is to minimize an outer stochastic objective function subject to the solution set of an inner stochastic convex optimization problem. Formally, the problem is defined as
$F_{\operatorname{opt}}:=\min_{x\in{\mathbf{R}}^{d}}\left\{F(x):={\mathbb{E}}[f(x,\theta)]:x\in X_{\operatorname{opt}}\right\},\ \text{where }X_{\operatorname{opt}}:=\operatorname*{arg\,min}_{z\in X}\left\{G(z):={\mathbb{E}}[g(z,\xi)]\right\}.$ (1)
Here $X\subset\mathbb{R}^{d}$ is a compact convex set, and both $F$ and $G$ are smooth with $G$ is additionally convex, making the optimization problem (1) a possibly nonconvex problem with convex domain. This framework has broad applicability, including hyper-parameter optimization [16, 17, 39], meta-learning [17, 42], reinforcement learning [32] and game theory [50]. In particular, problem (1) arises in learning applications where $G$ is the prediction error of a model $x$ on data, and $F$ is an auxiliary objective (e.g., regularization, or a validation data set). The dependence on data is through the expectation operators and the random variables $\theta,\xi$. The bilevel formulation provides a tuning-free alternative to the usual regularization approach where we optimize $\sigma F(x)+G(x)$ for a carefully tuned parameter $\sigma$.
Despite its wide applications in machine learning there are three primary challenges in solving (1). First, we do not have an explicit representation of the optimal set $X_{\operatorname{opt}}$ in general, preventing us from using some common operations in optimization such as projection onto or linear optimization over $X_{\operatorname{opt}}$. An alternative is to reformulate (1) using its value function formulation
@@ 1.1 Related works
Simple bilevel optimization problems with nonconvex outer objectives
This setting has been studied by Jiang et al. [29] in the deterministic case and by Cao et al. [4] in the stochastic case. These works propose conditional gradient methods that require performing linear optimization over intersections of the base domain $X$ with a halfspace. In contrast, our approach removes this requirement by relying solely on linear optimization over $X$. We note that we do not address the more general case with both nonconvex inner and outer objectives considered in [21, 24, 5]. In the remainder of this section, we will focus on simple bilevel problems with convex inner and outer objectives.
@@
Simple bilevel optimization extends Tikhonov regularization [52] beyond $\ell_{2}$ penalties by considering a blended objective $\sigma F(x)+G(x)$, where $\sigma>0$ modulates the trade-off between inner and outer objectives. As $\sigma\to 0$, the problem converges to a bilevel form that prioritizes the inner objective. Friedlander and Tseng [18] established fundamental conditions ensuring solution existence for $\sigma>0$, forming the basis for iterative regularization, where $\sigma$ is gradually reduced. This idea was first implemented by Cabot [3] in the unconstrained case $X={\mathbf{R}}^{d}$ using a positive decreasing sequence $\sigma_{t}$, and later extended to general convex constraints by Dutta and Pandit [14]. While these methods use proximal steps, which can be computationally expensive, Solodov [49] proposed a more efficient gradient-based alternative. Helou and Simões [23] further generalized the framework to non-smooth settings via an $\epsilon$-subgradient method. These works show asymptotic convergence but lack convergence rates. Amini and Yousefian [1] analyzed an iterative regularized projected gradient method and established a convergence rate of $O(t^{-(1/2-b)})$ for the inner objective, assuming non-smooth $F$ and $G$, compact $X$, and strong convexity of $f$. Kaushik and Yousefian [30] later removed the strong convexity assumption and proved rates of $O(t^{-b})$ and $O(t^{-(1/2-b)})$ for the inner and outer objectives, respectively, for any $b\in(0,1/2)$. Malitsky [35] studied an accelerated variant based on Tseng’s method [53], achieving an $o(t^{-1})$ rate for the inner problem. More recently, Merchav et al. [37] proposed an accelerated proximal scheme with a $O(\log(t+1)/t)$ rate under smoothness assumption and $O(t^{-2})$ and $O(1/t^{2-\gamma})$ for inner and outer objectives, respectively, under a Hölderian error bound condition with modulus $\gamma\in(1,2)$. Our work also follows an iterative regularization approach, but avoids expensive projections or proximal operations by leveraging linear optimization oracles.
@@
Several important applications have domains $X$ where projection-based operations are inefficient, yet linear optimization over $X$ is easier. To address the limitations of projection-based methods, recent work has focused on projection-free bilevel optimization via linear optimization oracles. Jiang et al. [29] approximated $X_{\operatorname{opt}}$ by replacing the constraint $G(x)\leq G_{\operatorname{opt}}$ with a linear approximation, which removed the need for Slater’s condition, and applying the conditional gradient method, refining the approximation at each step. They established $O(t^{-1})$ convergence rates for both objectives under smoothness and compactness assumptions in case $f$ is convex. Under non-convex setting, they claimed the convergence rates of $O(t^{-1/2})$ to the stationary point of the problem. However, their method requires a pre-specified tolerance parameter $\epsilon_{g}$, and only guarantees $(\epsilon_{g}/2)$-infeasibility. Doron and Shtern [12] proposed an alternative approach based on sublevel sets of the outer function $F$. Their method performs conditional gradient updates over sets of the form $X\times\{x:F(x)\leq\alpha\}$, with a surrogate $\hat{G}_{t}$ that is updated iteratively. They achieved rates of $O(t^{-1})$ for the inner and $O(t^{-1/2})$ for the outer objective under composite structure in $G$ and an error bound on $F$. While linear optimization over $X$ may be efficient, this is not always true for sets like $X\cap H$, where $H$ is a halfspace, or $\{F(x)\leq\alpha\}$ unless $F$ has a special structure. To address this, Giang-Tran et al. [20] introduced a conditional gradient-based iterative regularization scheme that only requires linear optimization over $X$, albeit with slower rates: $O(t^{-p})$ and $O(t^{-(1-p)})$ for the inner and outer problems, respectively, for any $p\in(0,1)$. We extend this framework to the stochastic setting, maintaining the same oracle-based reliance on linear optimization over $X$.
@@
Stochastic bilevel optimization remains less explored. Jalilzadeh et al. [27] developed an iterative regularization-based stochastic extragradient algorithm, requiring $O(1/\epsilon_{f}^{4})$ and $O(1/\epsilon_{g}^{4})$ stochastic gradient queries to achieve $\epsilon$-optimality for both objectives. Cao et al. [4] extended the projection-free framework of Jiang et al. [29] to the stochastic setting. Their methods achieve $\tilde{O}(t^{-1})$ convergence rate when the noise distributions have finite support, and $\tilde{O}(t^{-1/2})$ convergence rates under sub-Gaussian noise. However, their method requires linear optimization over intersections of $X$ with a halfspace. In contrast, our work relaxes this requirement by relying solely on linear optimization over the base set $X$.
@@
Several alternative approaches exist for simple bilevel optimization, including sequential averaging schemes [44, 46, 36], sublevel set-based methods [2, 6, 58, 21], accelerated algorithms [45, 54, 7, 6] and primal-dual strategies [47]. These methods are less related to our framework and do not address the stochastic setting. More general bilevel problems have seen considerable attention, particularly in Stackelberg game formulations where the upper and lower levels model leader-follower dynamics [19, 8, 31, 28, 10, 33, 56, 34]. Recent extensions include contextual bilevel optimization with exogenous variables [55, 25, 51] and problems with functional constraints at the lower level [30, 40]. Our focus in this paper, however, is restricted to the stochastic simple bilevel problems.
@@ 1.2 Contributions
We present two iterative regularization methods for solving problem (1) in both stochastic and finite-sum settings. The key idea is to introduce a decreasing regularization sequence $\{\sigma_{t}\}_{t\geq 1}$, and at each iteration $t$, perform a gradient-based update on the composite objective $\sigma_{t}F+G$. As $\sigma_{t}$ decreases, the algorithm gradually shifts focus from $F$ to $G$, thereby steering $x_{t}$ toward points in $X_{\operatorname{opt}}$. At the same time the $\sigma_{t}F$ term encourages $x_{t}$ towards a point in $X_{\operatorname{opt}}$ that minimizes $F$. In the stochastic formulation (1), computing exact gradients of the upper and lower objectives is computationally expensive. To address this, we employ sample-based gradient estimators. Specifically, we use the STOchastic Recursive Momentum (STORM) estimator [9] in the one-sample stochastic setting, and the Stochastic Path-Integrated Differential EstimatoR (SPIDER[38, 15] in the finite-sum setting.
@@
    One-sample stochastic setting: We develop a projection-free method for stochastic simple bilevel optimization that under convexity of $F$ on the base set $X$ achieves high-probability convergence rates of $O(t^{-1/4}\sqrt{\log(dt^{2}/\delta)})$ for the outer-level problem and $O(t^{-1/4}\sqrt{\log(d/\delta)})$ for the inner-level problem, with probability at least $1-\delta$. Without convexity of $F$, the method enjoys the non-asymptotic convergence rates of $O(t^{-\frac{1}{7}}\log({dt^{2}}/{\delta}))$ for both the outer- and inner-level objectives. Compared to the state-of-the-art approach by Cao et al. [4, Algorithm 1], which achieves the $O(t^{-1/2}\sqrt{\log(td/\delta)})$ rate under convex settings and $O(t^{-1/3}\sqrt{\log(td/\delta)})$ under nonconvex settings for both the outer and inner problems, our method offers several key advantages. First, it eliminates the need for optimization over intersections of the base domain with a halfspace, requiring only linear optimization over the original constraint set. Second, it removes the dependency on their pre-specified tolerance parameter $\epsilon_{g}$, and the associated computationally expensive initialization procedure to find an $\epsilon_{g}$-optimal starting point. Third, our algorithm enjoys anytime guarantees. Overall, the proposed algorithm is a simple, single-loop procedure that is significantly easier to implement than that of [4].
@@
    Finite-sum setting: For problems with $n$ component functions, we establish convergence rates of $O(t^{-1/2}\sqrt{\log(t^{2}/\delta)})$ (outer) and $O(t^{-1/2}\sqrt{\log(1/\delta)})$ (inner) with probability $1-\delta$. Without convexity of $F$, the method enjoys the non-asymptotic convergence rates of $O(t^{-\frac{1}{4}}\log({dt^{2}}/{\delta}))$ for both the outer- and inner-level objectives. Although Cao et al. [4, Algorithm 2] achieves improved $O(t^{-1}\log(t)\sqrt{\log(t/\delta)})$ rates under convex settings and $O(t^{-1/2}\sqrt{\log(t/\delta)})$ rates under nonconvex settings for both objectives, it again relies on more computationally demanding halfspace-intersection oracles, requires an additional initialization step, depends on a fixed stepsize, and does not offer anytime guarantees.
@@
    Numerical validation: We demonstrate the practical efficiency of our methods on over-parametrized regression and dictionary learning tasks, showing significant speedups over existing projection-based and projection-free approaches. The experiments validate both the convergence rates and computational advantages of our framework.
The supplementary material is organized in the appendices as follows. Appendix˜A establishes the preparatory lemmas used throughout our analysis. Appendices B and C contain the detailed proofs for the results presented in Section˜2 and Section˜3, respectively. Appendix˜D provides additional material, including analyses of the convergence rate coefficients as well as details of the numerical implementations in Section˜4.
2 Algorithm and convergence results for the one-sample stochastic setting
In this section, we introduce the Iteratively Regularized Stochastic Conditional Gradient (IR-SCG) method, summarized in Algorithm˜1. This approach is a conditional gradient method that leverages the STORM estimator to progressively reduce noise variance via momentum-based updates, assuming sub-Gaussian noise. While the use of the STORM estimator in projection-free methods was first explored in [60], our IR-SCG algorithm offers several key advantages. Notably, unlike the method of Cao et al. [4], IR-SCG is an anytime algorithm that does not require careful initialization, preset step sizes, or knowledge of the total number of iterations $T$ in advance.
Data: Parameters $\{\alpha_{t}\}_{t\geq 0}\subseteq[0,1]$$\{\sigma_{t}\}_{t\geq 0}\subseteq\mathbb{R}_{++}$.